In the mechanism shown in fig., the crank O2A makes 300




r.p.m in the conte clock wise direction as shown. Find




(i) Angular velocity of link O1A. (ii) Velocity of slider at B.

1
Expert's answer
2022-02-23T10:46:01-0500

"V_{OA}=oa=OA*\omega_{AB}"

"=60mm*200rpm"

"=60*10^{-3}*\frac{200}{60}=0.2m/s"

ba =5.6cm*0.1=0.56m/s

BA=8.4cm*2*10=168mm

(i) "\omega_{BA}=\frac{v_{BA}}{BA}"

"=\frac{0.56m/s}{0.168m}=3.33rps"

="200rpm"

(ii) "v_{BA}=0.56m\/s"